OpenStudy (anonymous):

Find the coordinates of the center of the given circle.(x - 5)2 + (y + 3)2 = 25

hartnn (hartnn):

for \(\Large (x-h)^2+(y-k)^2 = r^2\) the center is (h,k)

hartnn (hartnn):

so how about your equation ? \((x - 5)^2 + (y + 3)^2 = 25\) what are h and k values ?

OpenStudy (anonymous):

5 and 3

hartnn (hartnn):

actually when we compare y- k with y+3 we get k = -3 makes sense ?

hartnn (hartnn):

so your center is just (5,-3) !

hartnn (hartnn):

and for radius, if asked, r^2 = 25 gives r= radius = 5 :)

OpenStudy (anonymous):

thank you so much :-)

hartnn (hartnn):

welcome ^_^ since you're new here, \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \)